OpenStudy (anonymous):

Need help with Trig please i dont understand cos θ - cos θ sin2θ

OpenStudy (anonymous):

Need help with trig pleasse :)

OpenStudy (anonymous):

Is this supposed to be an equation? Or what do you need to do with it? The way this is written at the moment it's just a trigonometric expression.

OpenStudy (anonymous):

I need to simplfy the equation

OpenStudy (anonymous):

Hmm I see, it's a bit of a rough statement to me, because simplifying a trigonometric expression usually involves restrictions such as "express in terms of cosx only"

OpenStudy (anonymous):

well it says to use the identities but i dont get which one to use

OpenStudy (anonymous):

You may want to use \[\Large \sin(2x)=2\sin x \cos x \]

OpenStudy (anonymous):

so then i would sub in sin2x with 2sinx cosx?

OpenStudy (anonymous):

yes that's what I would do. But I doubt that this will simplify the equation a lot. It's one one possible way to express the trigonometric identity without a double angle.